The probability of occurrence of two events E and F are 0.25 and 0.50, respectively. the probability of their simultaneous occurrence is 0.14. the probability that neither E nor F occur is
\[ P(E \cup F) = P(E) + P(F) - P(E \cap F) \] \[ \Rightarrow P(E \cup F) = 0.25 + 0.50 - 0.14 = 0.61 \] Now, the probability that neither \(E\) nor \(F\) occurs is: \[ 1 - P(E \cup F) = 1 - 0.61 = 0.39 \]
Reinforce your conceptual grasp with related questions from NIMCET 2021:
Take timed full-length mock tests with authentic question distributions, negative marking (+12/-3 in Math), and instant All India Rank calculation.
Start Free Mock Test Series →